OpenStudy (anonymous):

A=P (1+r/n)nt Find A P=5,000 R=3% n=12 t=9

jhonyy9 (jhonyy9):

A= 5000*(1+(3/100)/12)*12*9 = 5000*(1+3/100 *1/12)*12*9 = =5000*(1+1/400)*108 =540000*(401/400)=5400*401/4 =1350*401= =541350 so A=541350